Solve by using the square root property.
2(3x-2)^2-10=0

Answer by lwsshak3(11628)

Solve by using the square root property.
2(3x-2)^2-10=0
2(3x-2)^2=10
(3x-2)^2=5
take sqrt of both sides
3x-2=±√5
3x=2±√5
x=(2±√5)/3